OpenStudy (anonymous):

Find the integral :

OpenStudy (anonymous):

\[\int\limits_{}^{}\left(1/x +2/x^2+ 3/x^3\right)\]

OpenStudy (amistre64):

convert it to negative exponents and integrate each term

OpenStudy (amistre64):

\[1 \int\limits_{} x^{-1} +2\int\limits_{}x^{-2} +3\int\limits_{}x^-3\]

OpenStudy (anonymous):

i have this for the answer \[\left| x \right| \ln 2/x +3/x^2 \]

OpenStudy (amistre64):

\[\ln(x) +\frac{2x^{(-1+1)}}{-1+1} + \frac{3x^{(-3+1)}}{-3+1}\]

OpenStudy (amistre64):

^(-2+1)...typoed it lol

OpenStudy (amistre64):

ln(x) -(2/x) -(3/2x^2)

OpenStudy (amistre64):

....+C ; cant forget the +C at the end

OpenStudy (anonymous):

ok thanks i missed so steps but i got it thank

OpenStudy (amistre64):

i missed some to, but I think I wrangled them all in at the end :)